# The Hidden Role of Anisotropies in Shaping Structure Formation in Cosmological N-Body Simulations

### v1

@@ V.2 Large box simulation
Figure 13: Heat-map of the ADPD for a slice of thickness $\Delta z=10$ Mpc of the large box CDM simulation at different redshifts: first row from right to left: $z=49,20,15,10$; second row $z=7.7,5.2,3.6,2.4$; third row $z=1.49,0.84,0.36,0$
@@
Figure 14: Point distribution for the same slices of Fig.13 (random sampling 10%).
@@
Figure 15: Angular variance of the ADPD for the large-box simulation at redshifts $z=50$ and $z=7$, shown alongside the corresponding result for a Poisson distribution with the same number of particles.
@@
Figure 16: Variance growth factor for the large-box simulation. Top panel: From redshift $z=20$ to $z=7$, the growth factor decreases as the imprint of the initial lattice is progressively erased, particles have moved, on average, by a distance comparable to the pre-initial lattice spacing. Bottom panel: After reaching a minimum at $z\approx 7$, the growth factor increases significantly, indicating the amplification of anisotropic structures seeded by residual features in the initial conditions. In this case the non-linear regime, occurs at very small scales, i.e. $r<10$ Mpc/$h$, compared to the simulation box.
[…]
Specifically, the ADPD heatmap at $z=49$ exhibits prominent vertical striping in angle $\theta$, with strong, repeating peaks at $\theta=0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, along with secondary features at $45^{\circ}$ intervals (see the upper left panel of Fig.3). These patterns directly reflect the inherent anisotropy of the cubic lattice, which features preferred directions aligned with the Cartesian axes and diagonals. Because the displacements are small, their contribution to the ADPD is negligible at this stage, the anisotropic lattice signature dominates.
@@
In contrast, at $z\approx 7$ the average particle displacement becomes comparable to the lattice spacing, effectively scrambling the initial grid configuration. The sharp lattice-induced striping observed in the ICs is largely washed out, indicating that the pairwise angular distribution is no longer governed by the pre-initial condition. Instead, the ADPD begins to reflect the geometry of the displacement field, including nascent features such as filaments and walls.
@@
Figure 17 shows a zoomed view of the 2D point distribution for the ICs of the large box at redshifts $z=49$ and $z=7.7$. In the early snapshot, the lattice structure remains almost unchanged due to the small amplitude of the displacement field at such high redshift. By contrast, in the later snapshot, the lattice imprint has been entirely erased by gravitational evolution. This transition is consistent with the behavior observed in the corresponding ADPD heat-maps.
@@
Figure 17: Left panel: zoom of the point distribution for the ICs of the large box at $z=49$ Right panel: zoom of the point distribution at $z=7.7$
@@
At $z\approx 7$, anisotropies seeded by the displacement field become dominant. Notably, the angular variance at this epoch exceeds that of a Poisson distribution with the same number of particles, confirming that the observed anisotropies are not statistical fluctuations, but systematic deviations induced by the coupling between an isotropic displacement field and a non-isotropic pre-initial lattice. Even in this case, we note that $\sigma^{2}_{\theta,\text{IC}}(r)\approx\text{const.}$, whereas the theoretical expectation scales as $\sim r^{-4}$ (see Eq. 27).
@@
These findings imply that even low-amplitude initial anisotropies can significantly influence the formation of large-scale structures, particularly in the linear and quasi-linear regimes where dynamics are still sensitive to ICs. While non-linear evolution at small scales eventually erases this memory, the early imprint of anisotropies remains relevant at scales $r>$ Mpc/h/h.
@@ VI Conclusions
These anisotropies are not merely transient or restricted to early times. Even when particles are displaced by more than the lattice spacing, the imprint of the initial configuration persists in the ADPD. This confirms that the anisotropic signature originates from the coupling of a structured (non-isotropic) pre-initial condition with an isotropic displacement field, producing a statistically anisotropic realization. As such, these features should be interpreted as numerical artifacts, discreteness effects not present in the theoretical model, but capable of influencing the morphology and clustering properties of the evolved matter distribution.
To demonstrate the impact of these artifacts, we compared two simulations: one with a small box ($L=62.5$ Mpc/h/h, $N=256^{3}$, $z_{\mathrm{IC}}=9$) where the displacement amplitude exceeds the lattice spacing, and another with a large box ($L=1000$ Mpc/h/h, $N=1024^{3}$, $z_{\mathrm{IC}}=49$) where it does not. In the latter case, the ADPD clearly retains the signature of the initial grid, indicating that spurious anisotropies persist when the particle initial displacement from the pre-initial position is insufficient to erase the lattice imprint.
@@
These results emphasize the limitations of lattice-based initial conditions in accurately modeling a statistically isotropic universe. They underscore the need for diagnostics sensitive to directional features — such as the ADPD — and motivate the adoption of alternative pre-initial configurations, particularly glass-like realizations. Glasses, constructed through reversed-sign gravitational relaxation, are designed to minimize artificial symmetries and suppress long-range correlations, making them a theoretically preferable choice for precision cosmological simulations. A detailed analysis of this case will be presented in a forthcoming work.
In conclusion, the hidden role of anisotropies introduced by pre-initial conditions is non-negligible. They represent a subtle but significant source of systematic error that can impact the fidelity of cosmological simulations. The ADPD provides a novel and robust method for identifying, characterizing, and ultimately mitigating these artifacts, paving the way for more accurate modeling of large-scale structure formation.
@@ Acknowledgements.
I wish to thank Benedikt Diemer for kindly providing the simulation data used in this study and for clarifying several aspects of the generation of initial conditions. I am also grateful to Marco Bruni, Roberto Capuzzo-Dolcetta, Andrea Gabrielli, Marco Galoppo and Michael Joyce for insightful discussions and valuable comments that helped improve the interpretation and presentation of the results.

### v2

@@ V.2 Large box simulation
Figure 13: Top panel: Point distribution in a zoomed region of the initial conditions for the large-box simulation at redshift $z_{\mathrm{IC}}=49$ (units: Mpc/$h$). Bottom panel: Heatmap of the Angular Distribution of Pairwise Distances for the same region of the ICs shown above. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
@@
Figure 14: Heat-map of the ADPD for a slice of thickness $\Delta z=10$ Mpc of the large box CDM simulation at different redshifts (the ICs was shown in Fig.13): first row from right to left: $z=49,20,15,10$; second row $z=7.7,5.2,3.6,2.4$; third row $z=1.49,0.84,0.36,0$. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
@@
Figure 15: Point distribution for the same slices of Fig.14 (random sampling 10%). Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
@@
Figure 16: Angular variance of the ADPD for the large-box simulation at redshifts $z_{\mathrm{IC}}=49$ and $z=7$, shown alongside the corresponding result for a Poisson distribution with the same number of particles. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
@@
Figure 17: Variance growth factor for the large-box simulation. Top panel: From redshift $z=20$ to $z=7$, the growth factor decreases as the imprint of the initial lattice is progressively erased, particles have moved, on average, by a distance comparable to the pre-initial lattice spacing. Bottom panel: After reaching a minimum at $z\approx 7$, the growth factor increases significantly, indicating the amplification of anisotropic structures seeded by residual features in the initial conditions. In this case the non-linear regime, occurs at very small scales, i.e. $r<10$ Mpc/$h$, compared to the simulation box. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
Fig. 13 displays the ADPD heatmap for the initial particle configuration at redshift $z_{\mathrm{IC}}=49$ measured in a slice of thickness $\Delta z=10$ Mpc/h/h. Similar to the case of the small-box simulation (see Fig. 7), the ADPD heatmap of the ICs exhibits patterns that directly reflect the intrinsic anisotropy of the underlying cubic lattice, characterized by preferred directions along the Cartesian axes and diagonals. Since the initial displacements are small, their contribution to the ADPD is negligible at this stage, and the anisotropic lattice signature remains dominant.
The redshift evolution of the ADPD measured in a slice of thickness $\Delta z=10$ Mpc/h/h is shown in Fig. 14 whereas the corresponding point distributions in these slices are shown in Fig. 15. The redshift slices in the range $49<z<10$ (first row of Fig. 14) allow us to track the progressive washing out of the initial lattice structure. This temporal evolution is clearly visible in the ADPD heatmaps of the large-box simulation, whereas for the small-box case this was not shown (see Fig. 9).
@@
At $z\approx 7$ the average particle displacement becomes comparable to the lattice spacing, effectively scrambling the initial grid configuration. The sharp lattice-induced striping observed in the ICs is largely washed out, indicating that the pairwise angular distribution is no longer governed by the pre-ICs. Instead, the ADPD begins to reflect the geometry of the displacement field, including nascent features such as filaments and walls.
@@
Fig. 18 shows a zoomed view of the two-dimensional point distribution in the large-box simulation at redshifts $z=20$ and $z=7.7$. In the early snapshot, the lattice structure—clearly visible in the ICs (see Fig. 13)—remains largely preserved, reflecting the small amplitude of the displacement field at such high redshift. By contrast, in the later snapshot, the imprint of the lattice has been completely erased by gravitational evolution. This transition is consistent with the behavior observed in the corresponding ADPD heatmaps.
@@
Figure 18: Left panel: zoom of the point distribution for the ICs of the large box at $z=20$. Right panel: zoom of the point distribution at $z=7.7$. Length scales are expressed in units of the average distance between nearest neighbors $\Lambda$.
@@
At $z\approx 7$, anisotropies seeded by the displacement field become dominant. Notably, the angular variance at this epoch exceeds that of a Poisson distribution with the same number of particles, confirming that the observed anisotropies are not statistical fluctuations, but systematic deviations induced by the coupling between an isotropic displacement field and a non-isotropic pre-initial lattice. Even in this case (see Fig.16), as for the small-box case (see Fig.12), we note that $\sigma^{2}_{\theta,\text{IC}}(r)\approx\text{const.}$, whereas the theoretical expectation scales as $\propto r^{-4}$ (see Eq. 27).
@@
These findings imply that even low-amplitude initial anisotropies can significantly influence the formation of large-scale structures, particularly in the linear and quasi-linear regimes where dynamics are still sensitive to ICs. While non-linear evolution at small scales eventually erases this memory, the early imprint of anisotropies remains relevant at scales $r>10$ Mpc/h/h.
@@ VI Conclusions
These anisotropies are not merely transient or restricted to early times. Even when particles are displaced by more than the lattice spacing, the imprint of the initial configuration persists in the ADPD. This confirms that the anisotropic signature originates from the coupling of a structured (non-isotropic) pre-ICs with an isotropic displacement field, producing a statistically anisotropic realization. As such, these features should be interpreted as numerical artifacts, discreteness effects not present in the theoretical model, but capable of influencing the morphology and clustering properties of the evolved matter distribution.
To demonstrate the impact of these artifacts, we compared two simulations: one with a small box ($L=62.5$ Mpc/h/h, $N=256^{3}$, $z_{\mathrm{IC}}=49$) and another with a large box ($L=1000$ Mpc/h/h, $N=1024^{3}$, $z_{\mathrm{IC}}=49$). In both cases, the amplitude of the initial displacement is smaller than the lattice spacing. As a result, the ADPD clearly retains the imprint of the underlying grid, indicating that spurious anisotropies persist when the particle displacements from their pre-initial positions are too small to effectively erase the lattice-induced anisotropy.
@@
These results highlight the limitations of lattice-based ICs in accurately modeling a statistically isotropic universe. They underscore the importance of employing diagnostics that are sensitive to directional features — such as the ADPD — and motivate the adoption of alternative pre-initial configurations, particularly glass-like realizations. Glasses, constructed through reversed-sign gravitational relaxation, are specifically designed to minimize artificial symmetries and suppress long-range correlations, making them preferable choice for pre-ICs of cosmological simulations.
Quantifying the anisotropy induced by different types of pre-ICs, and understanding its evolution during cosmic structure formation, is a highly relevant and timely question. This issue warrants a systematic investigation, especially in view of its potential impact on late-time observables beyond the two-point statistics — such as the bispectrum, measures of filamentarity, and topological descriptors.
A detailed analysis will be presented in a forthcoming work, which will involve the generation and analysis of simulations based on a variety of pre-initial configurations designed to reduce artificial symmetries. In particular, alternatives to the standard simple cubic (SC) lattice, include: the following possibilities. (i) Face-Centered Cubic (FCC): A highly symmetric and dense lattice, often regarded as the densest sphere packing in 3D, potentially providing a better approximation to isotropy than the SC lattice. (ii) Body-Centered Cubic (BCC): A configuration with an additional point at the center of each cube, offering higher packing efficiency than SC. (iii) Hexagonal Close-Packed (HCP): Composed of stacked hexagonal layers, this configuration is efficient in packing but exhibits reduced 3D symmetry. (iv) Tetrahedral and Octahedral Lattices: Based on repeating Platonic solids, these structures can introduce distinctive anisotropic signatures. (v) Quasi-crystalline Lattices: Aperiodic yet ordered structures that break translational invariance while maintaining long-range coherence. (vi) Voronoi and Delaunay Tessellations: Derived from regular or Poisson seed distributions, these methods are of particular interest in mesh-based modeling approaches. (vii) Glass-like Configurations: For instance, gravitational glasses, which are explicitly designed to be statistically isotropic and devoid of preferred directions.
Our main conclusion is that the role of anisotropies introduced by pre-ICs is non-negligible. They represent a subtle but significant source of systematic error that can impact the fidelity of cosmological simulations. The ADPD provides a novel and robust method for identifying, characterizing, and ultimately mitigating these artifacts, paving the way for more accurate modeling of large-scale structure formation.
@@ Acknowledgements.
I wish to thank Benedikt Diemer for kindly providing the simulation data used in this study and for clarifying several aspects of the generation of ICs. I am also grateful to Marco Bruni, Roberto Capuzzo-Dolcetta, Andrea Gabrielli, Marco Galoppo and Michael Joyce for insightful discussions and valuable comments that helped improve the interpretation and presentation of the results.